### v1

# Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young M4 + [T7 + T8] hierarchical triple

Abstract
Benchmark brown dwarfs in wide binary systems are crucial for characterizing substellar objects and calibrating atmospheric and evolutionary models. However, brown dwarf benchmarks with subsolar metallicity, very cool temperatures, or suitability for dynamical mass measurements are rare, limiting our understanding across the full range of mass, age, and metallicity. We present the discovery of two new multiple systems containing T dwarf companions, identified through a targeted search using CatWISE and Gaia databases. L 122-88 AB is a wide binary comprising a mildly metal-poor M2 dwarf and a T5 dwarf, separated by 215.6 arcsec at a distance of 33.1060.014 pc. Atmospheric model fitting to the near infrared spectrum of L 122-88 A suggests a mildly metal-poor composition ([Fe/H] = 0.2). UPM J10403551 AB is a candidate hierarchical triple system at 25.2830.013 pc, consisting of an M4 dwarf and a probable unresolved spectral binary of T7 and T8 dwarfs, separated by 65.48 arcsec from the primary. The H emission detected in UPM J10403551 A indicates an age range of 0.3-2.0 Gyr. This age estimate suggests that the T8 component has a mass between 9 and 28 Jupiter masses, potentially classifying it as a planetary-mass object. These systems augment the sample of benchmark brown dwarfs, particularly in the underexplored regime of cool temperature, providing valuable opportunities for refining our understanding of substellar objects.
keywords:
binaries: general – brown dwarfs – stars: late-type – stars: low-mass – gaseous planets1 Introduction
Brown dwarfs (BDs) are the most recently discovered major population on the Hertzsprung-Russell (H-R) diagram. They mostly form like stars but lack sufficient mass to sustain steady hydrogen fusion. Consequently, BDs represent the low-mass tail of the initial mass function (IMF) and are crucial for constraining it (Kirkpatrick et al., 2024). Bridging the gap between stars and planets, BDs share similar physical and atmospheric properties with gas giant planets (Hatzes & Rauer, 2015), providing essential references for the characterization of exoplanets.
To address the scientific objectives outlined above, samples of well-characterized BDs are required. However, BDs are challenging to characterize due to their faintness, cool temperatures, and the mass–age degeneracy inherent in their evolution. Without steady hydrogen burning, BDs cool over time, altering their spectral features and types (e.g. fig. 8 in Burrows et al., 2001). Therefore, benchmark BD samples with known mass, age, metallicity, and luminosity, spanning a wide parameter space, are needed to test atmospheric and evolutionary models. This need is especially pressing in the era of ongoing and upcoming deep surveys such as Euclid (Laureijs et al., 2011), the Roman Space Telescope (Spergel et al., 2015), the Rubin Observatory (LSST Science Collaboration et al., 2017), and the Chinese Space Station Telescope (CSST, Zhan, 2011).
BDs in nearby open clusters or moving groups with well-established ages serve as valuable age benchmarks (e.g., Teide 1, Rebolo et al. 1995; Luhman 16 AB, Luhman 2013). However, these clusters and groups are predominantly young ( 1 Gyr) and relatively distant ( 100 pc). In contrast, field BDs constitute the majority of the BD population, span a broader range of ages and metallicities, with many residing in the solar neighbourhood. Yet, direct measurements of age or mass for individual field BDs remain elusive. Nonetheless, BDs that are wide companions to well-characterized stars provide ideal benchmarks (e.g., Gliese 229 B, Nakajima et al. 1995; Indi B, Scholz et al. 2003; McCaughrean et al. 2004; King et al. 2010). Binary systems are expected to have coeval formation, sharing common age, metallicity, and distance—parameters more readily determined for stars than for BDs. This characteristic makes such systems particularly valuable for BD studies.
The T dwarf spectral class was established to classify BDs cool enough to exhibit methane absorption in their near-infrared (NIR) spectra (Burgasser et al., 2002; Geballe et al., 2002). T dwarfs have effective temperatures () between 500–1300 K (Kirkpatrick et al., 2021). They constitute the majority of field BDs (e.g. Zhang et al., 2019). However, T dwarfs are relatively rare in observations due to their cool temperatures and low luminosities. About 900 spectroscopically confirmed T dwarfs are known to date, mostly within 30 pc of the Sun (see summary in Best et al., 2024). Two-thirds of known T dwarfs have spectral types of T5–T9, owing to the rapid evolution of BDs through the L/T transition (e.g. Best et al., 2021).
Currently, there are about 30 T dwarfs known with stellar companions (summarised in Best et al., 2024), with significant contributions from the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010). To expand the parameter space of benchmark T dwarfs and improve their characterization, we searched for late-type T dwarfs with wide stellar companions using the CatWISE2020 catalog (Marocco et al., 2021) and the third data release (DR3) of Gaia (Gaia Collaboration et al., 2023). In this paper, we present the discovery of two new M + T dwarf wide binary systems. Section 2 describes the selection of wide binary candidates. Spectroscopic observation and characterization of these two systems are presented in Sections 3 and 4, respectively. A summary and conclusions are presented in Section 5.
2 Selection of wide binaries
Wide binaries ( 100 au) exhibit negligible orbital motion over decadal timescales due to their extremely long orbital periods ( 1 kyr). In the solar neighbourhood ( 100 pc), the proper motions (PMs) of wide binary components significantly exceed their orbital motions, resulting in apparent common PM over decades. This characteristic makes common PM an effective method for identifying nearby wide binaries (e.g., Zhang, 2019; Zhang et al., 2024).
T dwarfs emitting primarily in the infrared, only those in the solar neighbourhood are detectable by large-scale infrared surveys. The CatWISE database, providing mid-infrared photometry and PMs for nearby objects, serves as a powerful tool for identifying nearby T dwarf wide binaries.
We focused on late-type T dwarfs, which exhibit redder colour (Eq. 1) and suffer less contamination, e.g. from poorly measured background objects, than earlier types. Candidate selection from CatWISE employed the following criteria:
In our selection, we required signal to noise ratio (SNR) of better than 5.0 for W1 and W2 detections (Eq. 2). SNR of 5.0 corresponds to an error of 0.22 in W1 or W2 magnitudes. To mitigate contamination from background stars in the low-resolution (275 per pixel) unWISE images (Lang, 2014) used for CatWISE, we excluded the Galactic plane (Eq. 5). We select objects without identified artifacts in both W1 and W2 band unWISE images (Eq. 7) and with only one blend component used in the fit (Eq. 6). Approximately 9,540 objects in CatWISE met these criteria (Eqs 1–7).
We then searched for wide companions of these objects in the Gaia Catalogue of Nearby Stars (GCNS; Gaia Collaboration et al., 2021) using common PM. Our common PM pair criteria were:
Given the relatively large PM errors in CatWISE, we required significant PM (Eq. 3) but relaxed the PM constraint in the GCNS cross-match (Eq. 8). The distance limit of 40 pc (Eq. 9) corresponds to the -band detection limit of CatWISE for late T dwarfs (e.g. table 6 of Zhang et al., 2019). We limited the binary separation to within 10′ (corresponding to 24,000 au at 40 pc), beyond which the number of random contaminations arises, partially due to the large PM errors in CatWISE.
Our search yielded 15 candidate common PM pairs. After visual inspection using data from the Dark Energy Survey (DES; Abbott et al., 2021) and the DESI Legacy Imaging Surveys (Dey et al., 2019a), we excluded seven pairs of which the faint component has poor unWISE images, has Dark Energy Camera (DECam) deep -band non-detection (), or has no DECam observation and non-detection by other shallower surveys. The remaining eight T dwarf candidates, all with clear DECam -band detections, comprise two newly identified wide binaries and six known wide systems (Gliese 570 A & D, K4V + T7.5, Burgasser et al. 2000; G 204-39 AB, M3+T6.5, Faherty et al. 2010; SDSS J1416+13 AB, sdL7+sdT7.5, Scholz 2010; Burningham et al. 2010; HIP 73786 AB, sdM4+sdT5.5, Murray et al. 2011; Zhang et al. 2019; LHS 6176 AB, K8V+T8p, Burningham et al. 2013; L 34-26 + COCONUTS-2b, M3V+T9, Zhang et al. 2021). Fig. 1 shows the PMs of these common PM pairs.
The two new binaries, L 122-88 AB and UPM J10403551 AB have been independently reported recently. Marocco et al. (2024) classified UCAC3 52-1038 B (i.e. L 122-88 B) as a T6 dwarf based on photometric colours, and classified UCAC3 52-1038 A (i.e. L 122-88 A) as an M2 dwarf based on its low-resolution optical spectrum observed on 2021 December 25. Rothermich et al. (2024) classified UPM J1040–3551 A as an M3.8 dwarf based on photometric colours, and classified UPM J1040–3551 B as T7 dwarf based on a low resolution near infrared spectrum observed on 2023 June 30. Figure 2 shows the finding charts of these two new binaries. Their properties are listed in Table 1. The binary nature of these two newly identified pairs is corroborated by the concordance between the Gaia parallax-derived distances of the stellar components and the spectroscopically determined distances of their BD companions (Section 4.2).
L 122-88 AB and UPM J10403551 AB are located at distances of 33 and 25 pc respectively. They were missed by earlier searches with WISE and AllWISE (Wright et al., 2010; Mainzer et al., 2011), possibly because T dwarf candidates within 20 pc were prioritised in those spectroscopic follow up programs, to achieve a more complete volume-limited sample (e.g., Kirkpatrick et al., 2011, 2021; Mace et al., 2013). Another seven known T dwarf binaries were selected as T dwarf candidates by Eqs 1–7, but were not recovered as wide binaries mainly because of large uncertainties of their CatWISE PMs. We recovered most of them by double the limits in Eqs 3 and 4, but no new binary T dwarf candidate was found.
Primary | L 122-88 A | UPM J10403551 A | Ref. |
CWISE | J003124.30641354.9 | J104055.33355130.9 | |
Gaia DR3 | 4708465482876995712 | 5443160355149610368 | (1) |
(2016) | (1) | ||
(2016) | (1) | ||
Spectral type | M2 | M4 | (2) |
Age (Gyr) | 2.5-9.9 | 0.3-2.0 | (3,2) |
RUWE | 1.32 | 1.47 | (1) |
12.209 | 13.230 | (1) | |
13.346 | 14.863 | (1) | |
11.151 | 11.997 | (1) | |
(2MASS) | 9.812 0.026 | 10.302 0.023 | (4) |
(2MASS) | 9.248 0.023 | 9.676 0.022 | (4) |
(2MASS) | 8.991 0.019 | 9.446 0.023 | (4) |
W1 | 8.868 0.012 | 9.280 0.012 | (5) |
W2 | 8.729 0.007 | 9.087 0.007 | (5) |
(mas) | 30.206 0.013 | 39.552 0.020 | (1) |
Distance (pc) | 33.106 0.014 | 25.283 0.013 | (1) |
(mas yr-1) | 227.556 0.014 | 128.580 0.019 | (1) |
(mas yr-1) | 269.121 0.014 | 10.646 0.018 | (1) |
(km s-1) | (1) | ||
RV (km s-1) | 15.63 0.31 | 0.09 3.72 | (1) |
6.391 0.020 | 7.432 0.024 | (2) | |
(K) | 3618 157 | 3244 157 | (6) |
(K) | 3500 50 | 3200 50 | (2) |
Mass (M⊙) | 0.378 0.020 | 0.235 0.020 | (6) |
Companion B | L 122-88 B | UPM J10403551 Ba/Bb | |
CWISE | J003147.79641123.1 | J104053.42355029.7 | |
(WISE) | (3) | ||
(WISE) | (3) | ||
Mean mjd (day) | 57265.87 | 57124.73 | (3) |
Spectral type | d/sdT5 | T7 + T8 | (2) |
(DECaPS) | 25.149 0.427 | — | (7) |
(DECaPS) | 23.824 0.216 | — | (7) |
(DECaPS) | 21.107 0.035 | 21.133 | (7) |
(DECaPS) | 19.821 0.046 | — | (7) |
(VHS) | 17.601 0.015 | 17.393 0.022 | (8) |
(VHS) | 18.417 0.239 | 17.344 0.128 | (8) |
W1 | 17.380 0.055 | 16.773 0.041 | (5) |
W2 | 15.351 0.029 | 14.940 0.025 | (5) |
Distance (pc) | 45 | a20, b29 | (2) |
(mas yr-1) | 222 33 | 135 27 | (5) |
(mas yr-1) | 293 30 | 6 28 | (5) |
Mass (M⊙) | 0.05–0.067 | 0.011–0.033/0.009–0.027 | (2) |
Separation (″) | 215.6 | 65.48 | (2) |
Proj. sep. (au) | 7138.4 0.5 | 1655.6 0.3 | (2) |
cProj. sep. () | 0.034 | 0.009 | (2) |
(J) | (2) |
aAverage spectroscopic distance as a single T7 dwarf based on spectral type – absolute magnitudes in bands (Zhang et al., 2019).
bAverage spectroscopic distance as a T7+T8 binary. The extracted apparent magnitude of the T7 companion is used in the calculation.
cProjected separation in Jacobi radius (, see Section 4.3.)
Name | SpT | UT date | Telescope | Instrument | Slit | Seeing | Airm | Wavelength | Resolution | |
(arcsec) | (arcsec) | (nm) | () | (sec) | ||||||
UPM J1040–3551 A | M5 | 2023-12-29 | SOAR | Goodman/400 | 1.0 | 1.45 | 1.04 | 492-897 | 2000 | |
UPM J1040–3551 B | T7+T8 | 2024-03-26 | SOAR | TripleSpec | 1.1 | 0.91 | 1.07 | 940-2470 | 3500 | |
L 122-88 A | M2 | 2023-12-29 | SOAR | Goodman/400 | 1.0 | 1.55 | 1.89 | 492-897 | 2000 | |
[…]
L 122-88 B | d/sdT5 | 2024-07-11 | SOAR | TripleSpec | 1.1 | 1.24 | 1.35 | 940-2470 | 3500 | |
[…]
3 Spectroscopic observations
We conducted spectroscopic follow-up observations of L 122-88 AB and UPM J10403551 AB prior to their independent reporting. As spectra for L 122-88 B and UPM J10403551 A were not available in the literature, we continued our spectroscopic campaign to complete the analysis of these new binary systems. Table 2 summarises the details of our observations.
3.1 SOAR/Goodman
L 122-88 A and UPM J1040–3551 A were observed using the Goodman High Throughput Spectrograph (Goodman; Clemens et al., 2004) at the Southern Astrophysical Research Telescope (SOAR) in long-slit mode on 2023 December 29 (see Fig. 3). We used a slit and the 400 lines mm-1 grism in M2 mode, providing wavelength coverage from 492 to 897 nm with a resolving power of 2,000. Observations were conducted at airmasses of 1.89 and seeing of 1.55 for L 122-88 A, and airmasses of 1.04 and seeing of 1.45 for UPM J1040–3551 A.
Data reduction was performed using the Goodman HST Pipeline111https://soardocs.readthedocs.io/projects/goodman-pipeline/ for basic CCD reduction, including bias subtraction, flat-field correction, and cosmic ray removal. Further spectroscopic processing utilized tasks from the NOAO package in IRAF, including wavelength calibration, spectral extraction, telluric correction, and flux calibration. HIP 45167 served as the telluric standard star for flux calibration.
3.2 SOAR/TripleSpec
UPM J10403551 B and L 122-88 B were observed with the TripleSpec4.1 NIR Imaging Spectrograph (TripleSpec; Schlawin et al., 2014) at SOAR. TripleSpec provides simultaneous spectral coverage from 940 to 2470 nm (R3,500), spanning the entire photometric range (see Fig. 4). UPM J10403551 B was observed on 2024 March 26 (seeing: 091, airmass: 1.07), while L 122-88 B was observed on 2024 July 11 (seeing: 124, airmass: 1.35). We adopted the default ABBA dither pattern with eight 225-s exposures for UPM J1040–3551 B and twenty 200-s exposures for L 122-88 B.
Data reduction employed a modified version of the IDL-based Spextool pipeline (Cushing et al., 2004), adapted for SOAR. The process included flat-field correction, wavelength calibration using a CuHeAr arc lamp, removal of emission sky features by subtracting A and B exposures, and one-dimensional spectral extraction. Telluric correction and flux calibration were performed using observations of the standard star HIP54311. The resulting NIR spectra exhibit peak signal-to-noise ratios (SNR) of 10-15 at the centre of the -bands.
3.3 VLT/X-shooter
The X-shooter spectrograph (Vernet et al., 2011) at the Very Large Telescope comprises three spectroscopic arms (UVB, VIS, and NIR), each optimized for its spectral range (300-2500 nm). The optical to NIR spectrum of L 122-88 A was obtained using the X-shooter on 2024 June 16 (seeing: 15, airmass: 1.8). Figure 5 presents the combined, smoothed X-shooter spectrum of L 122-88 A.
L 122-88 B was observed with X-shooter on 2024 July 7 (seeing: 358) and 2024 July 13 (seeing: 129). Figure 6 shows the X-shooter spectrum of L 122-88 B, noting that the optical spectrum is contaminated by a nearby galaxy, in fact, we note the presence of a non-zero continuum in the 400-800 nm spectral region in which T dwarfs should barely have emit any flux (e.g., Burgasser et al., 2003b). The nearby galaxy is faint in the -band (see Fig. 2) but brighter than L 122-88 B in the optical. With a PM of 0368 yr-1 towards the north-east, L 122-88 B has moved to the top of the galaxy in 2024 when the X-shooter spectrum was taken. Meanwhile, the flux of the X-shooter spectrum beyond 2250 nm is poorly extracted/calibrated due to low SNR. These spectra were reduced using the ESO Reflex pipeline (Freudling et al., 2013).
4 Characterization
4.1 M dwarf companions
The Gaia Renormalized Unit Weight Error (RUWE) values for L 122-88 A and UPM J10403551 A (1.32 and 1.47, respectively) suggest these objects may potentially be unresolved binaries. However, their optical spectra appear consistent with normal M dwarfs (Fig. 3), indicating that if they are indeed binaries, each component likely possesses a similar spectral classification. Under such circumstances, given their identical distances, metallicity constraints derived from model fitting, whether applied to their composite spectra or to individual components, would yield consistent results. In this section, we examine their spectra as single objects.
4.1.1 L 122-88 A
We compared the optical spectra (Fig. 3) of L 122-88 A to those M0-M9 dwarf templates from Bochanski et al. (2007), and found it closely matches an M2 dwarf template. However, it exhibits slightly weaker TiO absorption-bands around 710 nm, which indicates a mildly metal-poor composition (e.g., Lépine et al., 2007; Jao et al., 2008; Zhang, 2019).
Analysis of Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) data using the TESSILATOR software (Binks & Günther, 2024) reveals a rotation signal of 2.4750.013 days in sector 2 data. Applying gyrochronology to this rotation period using GPgyro (Lu et al., 2024) yields an age estimate of 0.0290.010 Gyr. However, there is no evidence for this signal in other TESS sectors 28, 29, 68, and 69. Examination of the TESS data sector 2 release notes222https://ntrs.nasa.gov/citations/20190001284 for TIC 271503151 (L 122-88 A) does not suggest any particular issue with the sector 2 observation but for a V=13 object we would expect the signal to be visible in other sectors for a reliable detection. Marocco et al. (2024) suggested a significantly older kinematic age of 2.5-9.9 Gyr for L 122-88 A.
4.1.2 UPM J10403551 A
We compared the optical spectrum (Fig. 3) of UPM J10403551 A to those M0-M9 dwarf templates from Bochanski et al. (2007), and found it closely matches that of an M4 dwarf template. The presence of H emission in its spectrum indicates that it is an active M4 dwarf, which is also suggestive of a relatively young age. We measured the H equivalent width (H EW) of UPM J10403551 A to be between 2.33 and 3.78 Å. Applying the H activity-age correlation for M3-6 dwarfs from Kiman et al. (2021) (their fig. 6), this H EW range corresponds to an estimated age of 0.3-2.0 Gyr. The TESS data for UPM J10403551 A shows no indication of activity or any periodic signal.
4.1.3 Atmospheric properties
The optical spectrum is more sensitive to of M (sub)dwarfs, while the NIR spectrum is more sensitive to [Fe/H]. Therefore, the full optical-NIR spectral profile of M subdwarfs can have better constraints of their and [Fe/H]. To derive the atmospheric parameters of L 122-88 A, we fitted its X-shooter spectrum (300-2500 nm) to BT-Settl models with 2600 K K, [Fe/H] 0.5, and log = 5.0 (Allard, 2014). We used a typical log = 5.0 for M2-M4 dwarfs as log has relatively small affection on their spectra compared to and [Fe/H]. We applied smoothing to the X-shooter spectrum to enhance the SNR, though this degraded the resolving power to 1000. The BT-Settl model spectra were similarly degraded to match the resolution of the processed X-shooter spectrum.
The fitting procedure involved comparing the observed spectrum with a grid of models, varying in steps of 50 K and [Fe/H] in steps of 0.1 dex. We employed a least-squares method for the comparison. The best-fitting model for L 122-88 A yielded = 3500 50 K, [Fe/H] = , and log = 5.0 (Fig. 5), confirming its mildly metal-poor nature.
We applied a similar fitting procedure to the optical spectrum of UPM J10403551 A, using the same set of BT-Settl models. The best-fitting model for UPM J10403551 A resulted in = 3200 K, [Fe/H] = 0.0, and log = 5.0 (Fig. 3), indicating a solar metallicity for this object.
4.2 T dwarf companions
Figure 4 presents the NIR spectra of L 122-88 B and UPM J10403551 B obtained with TripleSpec. These spectra have been smoothed to facilitate comparison with T dwarf standards. Our analysis of L 122-88 B primarily relies on its TripleSpec spectrum, which offers higher SNR compared to its X-shooter spectrum (Fig. 6).
4.2.1 L 122-88 B
The NIR spectrum of L 122-88 B (Figs 4 and 6) fits well to that of the T5 standard 2MASS J15031961+2525196 (J1503+25; Burgasser et al., 2003a, 2004), but exhibits suppressed -band flux which is a subsolar metallicity feature of BDs (Burgasser et al., 2002; Zhang et al., 2017a). It bears a strong resemblance to the sdT5.5 subdwarf benchmark HIP 73786 B (Murray et al., 2011; Zhang et al., 2019), which has a metallicity of [Fe/H] inferred from its primary HIP 73786 A (Carollo et al., 2007; Soubiran et al., 2016; Arentsen et al., 2019). The slightly weaker -band flux suppression in L 122-88 B compared to HIP 73786 B suggests a marginally higher metallicity, consistent with the metallicity derived from model fitting of L 122-88 A ([Fe/H] = , Section 4.1). Consequently, we classify L 122-88 B as a mildly metal-poor T5 dwarf (d/sdT5).
HIP 73786 B exhibits a weak flux excess on the left-hand side of the -band peak (980-1080 nm), which is possibly caused by subsolar metallicity. This feature is also visible in spectra of other T subdwarfs with [Fe/H] and becomes more significant at lower metallicity (e.g., Pinfield et al., 2012; Burningham et al., 2014). The absence of the -band flux excess feature in L 122-88 B indicates a metallicity higher than that of T subdwarfs, which usually have [Fe/H] (e.g., fig. 9, Zhang et al., 2017b). This is still consistent with the metallicity derived from L 122-88 A.
To further investigate the mildly metal-poor nature of L 122-88 B, we compared its absolute magnitudes in the and -bands, computed using the Gaia parallax of the primary, to those of known T dwarfs and subdwarfs (Fig. 7). L 122-88 B exhibits slightly fainter and compared to typical T5 dwarfs, while its remains similar. This pattern is consistent with the behaviour of T5.5-9 subdwarfs, which show fainter and -band absolute magnitudes than their solar-metallicity counterparts, but indistinguishable -band absolute magnitudes (Zhang et al., 2019). Without considering its mildly metal-poor nature, L 122-88 B might be misclassified as a T6.51 dwarf based on spectral type versus absolute magnitude correlations (Fig. 7). Indeed, Marocco et al. (2024) classified it as a T6 dwarf based on spectral type versus colour correlations from Kirkpatrick et al. (2021).
We also estimated the distance of L 122-88 B by the spectral type – absolute magnitude correlations (fig. 10, Zhang et al., 2019). The average distance estimated from its magnitudes is 45 pc, which is consistent to the Gaia distance of L 122-88 A (33.1060.014) within the uncertainty. Of course, L 122-88 B exhibits a over estimated spectral distance due to its mildly metal-poor composition. This metallicity deficiency causes its actual absolute magnitudes to be fainter than the average values for T5 dwarfs used in the calculation.
4.2.2 UPM J10403551 B: a probable unresolved binary
The NIR spectrum of UPM J10403551 B closely resembles that of the T7 dwarf standard 2MASS J0727182+171001 (J0727+17; Burgasser et al., 2002, 2006). However, Fig. 4 reveals flux excesses in the and bands relative to the T7 standard when normalized at the -band peak. When rescaling the spectrum of UPM J1040-3551 B to better align with the T7 standard across , , and bands, we observe a 13% flux suppression in the band, a phenomenon neither previously observed nor explained in any single BD.
Several features of evidence suggest that UPM J10403551 B is not a T subdwarf: (1) it exhibits -band flux excess rather than the suppression typical of late-type T subdwarfs, (2) it has a relatively young age of 0.3-2.0 Gyr inferred from its primary (Section 4.1.2), and (3) it has solar metallicity as indicated by its primary’s spectrum (Fig. 3). While the -band flux excess could be attributed to its young age (e.g., Burningham et al., 2011), this explanation fails to account for the observed and band flux excesses.
Unresolved binary BDs may show unusual spectral features and colours, and flux excesses if have been treated as single objects (Burgasser et al., 2010; Zhang et al., 2010; Marocco et al., 2015; Zhang et al., 2018b; Kirkpatrick et al., 2021). Figure 7 demonstrates that UPM J10403551 B has significantly brighter and band absolute magnitudes than typical T7 dwarfs. Such over-luminosity in field ultracool dwarfs often indicates the presence of an unresolved companion.
To test this hypothesis, we conducted a binary spectral fitting analysis using single spectral templates in the Spex Prism Library Analysis Toolkit (Burgasser & Splat Development Team, 2017, SPLAT;). We selected 131 T4–T9 dwarfs with low-resolution ( 150) NIR spectra previously obtained with the SpeX spectrograph (Rayner et al., 2003) on the 3m NASA Infrared Telescope Facility. These templates exclude known or suspected binaries and spectra with very low S/N 15. Spectra were interpolated and resampled onto a common wavelength scale spanning 0.95–2.5 m at a constant resolution = 200, classified by comparison to NIR spectral standards (Burgasser et al., 2006), and then flux-calibrated using the Kirkpatrick et al. (2021) absolute MKO -band magnitude/spectral type relation. We then co-added pairs of spectra to create 6,715 unique binary templates. We compared both single and binary templates to the equivalently interpolated SOAR spectrum of J10403551 B following procedures described in (Burgasser et al., 2010), masking out regions of strong telluric absorption between 1.35–1.42 m and 1.80–1.95 m. Figure 8 compares the best fit single template from this sample, the T7 dwarf WISE J004024.88+090054.8 (Mace et al., 2013, J0040+0900;) to the best-fit binary template composed of J0040+0900 and the T8 dwarf ULAS J102940.52+093514.6 (Burningham et al., 2013; Thompson et al., 2013, J1029+0935;). The latter is a superior fit, notably providing a better reproduction of the , , and -band peaks. The difference in reduced between the single fit (10.6) and the binary fit (7.6) implies a 2% likelihood of equivalence based on the F-test statistic. However, we caution that other physical parameters, such as youth/low surface gravity, could also explain the modest deviations in the single template fit and may not be fully accounted for in the existing SPLAT template sample. The spectral fitting therefore provides supporting but not conclusive evidence of unresolved multiplicity.
To further investigate this binary hypothesis, we decomposed the absolute magnitudes of UPM J10403551 B into T7 and T8 components. Using the spectral type versus absolute magnitude correlations for L and T dwarfs (Fig. 7), we derived , , and differences between T7 and T8 dwarfs of 1.258, 1.248, 0.907, and 0.462 mag, respectively. The decomposed , , and values demonstrate improved alignment with these correlations, though they remain slightly brighter than average—contrary to the dimmer characteristics of late T subdwarfs. UPM J10403551 B appears underluminous as an unresolved binary only in the -band, which may be attributed to its young age. For instance, the young T5.5 dwarf SDSSp J111010.01+011613.1 (Geballe et al., 2002) is relatively faint in the -band, with a colour 0.34 mag bluer than the average for field T5.5 dwarfs (fig. 8i of Primeval VI, ).
The independent spectroscopic distance of UPM J10403551 B as a single T7 dwarf is 20 pc, which is 21% shorter than the Gaia distance of the M4 primary (25.2830.013 pc). We also decomposed the band apparent magnitudes of UPM J10403551 B as T7 + T8 close binary. The average spectroscopic distance of UPM J10403551 Ba as a T7 dwarf is 21 pc, which is 17% smaller than its UPM J10403551 A but consistent within the uncertainty. The spectral distance of UPM J10403551 B is under estimated due to its young age, which allows it to have brighter absolute magnitudes than the average values for T7 dwarfs used in the calculation.
Both the spectral fitting and the decomposition of absolute magnitudes strongly suggest that UPM J10403551 B is likely an unresolved T7 + T8 binary (UPM J10403551 Ba & Bb). UPM J10403551 ABab might be the third known triple system comprising a close binary of T dwarfs with a wide stellar companion, similar to Indi ABab (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010) or Gliese 229 ABab (Xuan et al., 2024).
UPM J10403551 B is detected but not resolved in the Visible and Infrared Survey Telescope for Astronomy (VISTA) Hemisphere Survey (VHS; McMahon et al., 2013) and the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019b). The VHS and DECaLS images have pixel sizes of 034 and 026, respectively. This allows us to place upper limits on the orbital semi-major axis ( au) and period ( yr) of the system. At a distance of 25 pc, the components can potentially can be resolved at 1.15 m by the Near-Infrared Camera (NIRCam) of the James Webb Space Telescope (JWST), provided the orbital period exceeds six years.
4.3 Binary properties
To estimate the masses of the T dwarf companions, we employed the evolutionary tracks of from the Sonora-Bobcat models (Marley et al., 2021), using their estimated ages and the average of T dwarfs with corresponding spectral types. As illustrated in Fig. 9, L 122-88 B has an estimated mass of 0.05–0.067 M for an age of 2.5-9.9 Gyr, while UPM J10403551 Ba and Bb have estimated masses of 0.012–0.04 M and 0.007–0.02 M⊙, respectively, for an age of 0.3-2.0 Gyr. Using the derived distances, projected separations, and masses of all components, we calculated the binding energies of L 122-88 AB and UPM J10403551 ABab to be J and J, respectively.
To assess the stability of these wide binary systems, we considered the Jacobi (or tidal) radius (), which represents the maximum separation for stable wide binary systems. The delineates the boundary beyond which the tidal field dominates over the gravitational attraction between the binary components. For wide binaries in the solar neighbourhood, can be calculated using equation (43) from Jiang & Tremaine (2010).
We found that L 122-88 AB and UPM J10403551 ABab have values of 1.02 pc and 0.87 pc, respectively. Their projected separations correspond to 0.034 and 0.009 , respectively, which are far smaller than their respective values. This analysis strongly suggests that both systems are gravitationally stable. The spectroscopic distances of L 122-88 B and UPM J10403551 B are also consistent to Gaia parallax distances of their primaries (see Table 1).
5 Summary and conclusions
We have conducted a search for late T dwarf wide binaries using CatWISE and Gaia DR3, resulting in the discovery of two new M + T dwarf wide binary systems. Both systems are confirmed to be gravitationally bound and stable based on their tidal radii. Optical and NIR spectroscopic follow-up observations were performed for all companions, enabling spectral classification and characterization.
The first system, L 122-88 AB, exhibits spectral features indicative of mild metal-poor composition, including a weaker 710 nm TiO absorption band in L 122-88 A and suppressed K-band flux in L 122-88 B. The best-fitting BT-Settl model to the 0.3-2.5 m spectrum of L 122-88 A yields = 3500 K, [Fe/H] = , and log = 5.0. We classify L 122-88 AB as a mildly metal-poor M2 + T5 wide binary.
The second system, UPM J10403551 AB, is likely a hierarchical triple system consisting of an unresolved T7 + T8 binary and an M4 dwarf wide companion at a distance of 25 pc from the Sun. While UPM J10403551 Ba/Bb are unresolved in survey images, their NIR spectrum is best fitted by a combined T7 + T8 template, explaining the observed excess luminosity. This system might be the third known triple system similar to Indi ABab or Gliese 229 ABab. Based on the H emission of the M4 primary, we estimate the system age to be 0.3-2 Gyr. At this relatively young age, the T8 component UPM J10403551 Bb has a mass between 9 and 28 Jupiter masses and potentially is a planetary mass object.
These newly discovered systems, L 122-88 AB and UPM J10403551 AB, augment the sample of benchmark BDs, particularly in the less-explored regimes of cool temperature. UPM J10403551 B potentially is a good target for dynamical mass measurement with high resolution imaging. They present valuable opportunities for refining atmospheric and evolutionary models, thereby enhancing our understanding of substellar objects approaching the planetary mass limit. Furthermore, these systems may provide crucial insights into bridging the gap between BDs and giant exoplanets, contributing to a more comprehensive understanding of the low-mass end of the stellar-substellar-planetary continuum.
Acknowledgements
Based on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). Based on observations collected at the European Southern Observatory under ESO programme 113.26V7.001.
This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Wide field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Lab- oratory/California Institute of Technology, funded by the National Aeronautics and Space Administration This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. SPLAT is an collaborative project of research students in the UCSD Cool Star Lab, aimed at developing research through the building of spectral analysis tools. Contributors to SPLAT have included Christian Aganze, Jessica Birky, Daniella Bardalez Gagliuffi, Adam Burgasser (PI), Caleb Choban, Andrew Davis, Ivanna Escala, Joshua Hazlett, Carolina Herrara Hernandez, Elizabeth Moreno Hilario, Aishwarya Iyer, Yuhui Jin, Mike Lopez, Dorsa Majidi, Diego Octavio Talavera Maya, Alex Mendez, Gretel Mercado, Niana Mohammed, Johnny Parra, Maitrayee Sahi, Adrian Suarez, Melisa Tallis, Tomoki Tamiya, Chris Theissen, and Russell van Linge.
Z.H.Z acknowledges the supports from the Jiangsu Province fundamental research programme (BK20211143), the Program for Innovative Talents and Entrepreneur in Jiangsu (JSSCTD202139), and the science research grants from the China Manned Space Project with NO. CMS-CSST-2021-A08. The work of F.N. is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. M.C.G.O. acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF “A way of making Europe” through project PID2022-137241NB-C42. N.L. acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (AEI-MCINN) under grant PID2022-137241NB-C41. B.G. acknowledges support from the Polish National Science Center (NCN) under SONATA grant No. 2021/43/D/ST9/0194. P.C. acknowledges financial support from the Spanish Virtual Observatory project funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 through grant PID2020-112949GB-I00. The authors would like to thank the anonymous reviewer for his/her valuable comments.
Data Availability
The data underlying this article are available in the article and in its online supplementary material.

### v3

@@
-Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young M4 + [T7 + T8] hierarchical triple
+Benchmark brown dwarfs – I. A blue M2 + T5 wide binary and a probable young [M4 + M4] + [T7 + T8] hierarchical quadruple
 Abstract
@@
 However, brown dwarf benchmarks with subsolar metallicity, very cool temperatures, or suitability for dynamical mass measurements are rare, limiting our understanding across the full range of mass, age, and metallicity.
-We present the discovery of two new multiple systems containing T dwarf companions, identified through a targeted search using CatWISE and Gaia databases.
+We present the discovery of two new multiple systems containing T dwarf companions, identified through a targeted search using CatWISE2020 and Gaia catalogues.
 L 122-88 AB is a wide binary comprising a mildly metal-poor M2 dwarf and a T5 dwarf, separated by 215.6 arcsec at a distance of 33.1060.014 pc.
 Atmospheric model fitting to the near infrared spectrum of L 122-88 A suggests a mildly metal-poor composition ([Fe/H] = 0.2).
-UPM J10403551 AB is a candidate hierarchical triple system at 25.2830.013 pc, consisting of an M4 dwarf and a probable unresolved spectral binary of T7 and T8 dwarfs, separated by 65.48 arcsec from the primary.
+UPM J10403551 AB is a candidate hierarchical quadruple system at 25.2830.013 pc, consisting of a likely astrometric binary of two M4 dwarfs and a probable unresolved spectral binary of T7 and T8 dwarfs, separated by 65.48 arcsec from the primary.
 The H emission detected in UPM J10403551 A indicates an age range of 0.3-2.0 Gyr.
@@
 keywords:
-binaries: general – brown dwarfs – stars: late-type – stars: low-mass – gaseous planets1 Introduction
-Brown dwarfs (BDs) are the most recently discovered major population on the Hertzsprung-Russell (H-R) diagram.
+planets and satellites: gaseous planets – binaries: general – brown dwarfs – stars: late-type – stars: low-mass1 Introduction
+Brown dwarfs (BDs) are the most recently discovered major population on the Hertzsprung-Russell diagram.
 They mostly form like stars but lack sufficient mass to sustain steady hydrogen fusion.
-Consequently, BDs represent the low-mass tail of the initial mass function (IMF) and are crucial for constraining it (Kirkpatrick et al., 2024).
+Consequently, BDs represent the low-mass tail of the initial mass function and are crucial for constraining it (Kirkpatrick et al., 2024).
 Bridging the gap between stars and planets, BDs share similar physical and atmospheric properties with gas giant planets (Hatzes & Rauer, 2015), providing essential references for the characterization of exoplanets.
@@
 fig.
-8 in Burrows et al., 2001).
+8 of Burrows et al., 2001).
 Therefore, benchmark BD samples with known mass, age, metallicity, and luminosity, spanning a wide parameter space, are needed to test atmospheric and evolutionary models.
-This need is especially pressing in the era of ongoing and upcoming deep surveys such as Euclid (Laureijs et al., 2011), the Roman Space Telescope (Spergel et al., 2015), the Rubin Observatory (LSST Science Collaboration et al., 2017), and the Chinese Space Station Telescope (CSST, Zhan, 2011).
-BDs in nearby open clusters or moving groups with well-established ages serve as valuable age benchmarks (e.g., Teide 1, Rebolo et al.
-1995; Luhman 16 AB, Luhman 2013).
+This need is especially pressing in the era of ongoing and upcoming deep surveys such as Euclid (Laureijs et al., 2011), the Roman Space Telescope (Spergel et al., 2015), the Rubin Observatory (LSST Science Collaboration et al., 2017), and the Chinese Space Station Telescope (CSST; Zhan, 2011).
+BDs in nearby open clusters or moving groups with well-established ages serve as valuable age benchmarks [e.g., Teide 1, (Rebolo et al., 1995) and Luhman 16 AB, (Luhman, 2013)].
 However, these clusters and groups are predominantly young ( 1 Gyr) and relatively distant ( 100 pc).
-In contrast, field BDs constitute the majority of the BD population, span a broader range of ages and metallicities, with many residing in the solar neighbourhood.
+In contrast, field BDs constitute the majority of the BD population, and span a broader range of ages and metallicities, with many residing in the solar neighbourhood.
 Yet, direct measurements of age or mass for individual field BDs remain elusive.
-Nonetheless, BDs that are wide companions to well-characterized stars provide ideal benchmarks (e.g., Gliese 229 B, Nakajima et al.
-1995; Indi B, Scholz et al.
-2003; McCaughrean et al.
-2004; King et al.
-2010).
-Binary systems are expected to have coeval formation, sharing common age, metallicity, and distance—parameters more readily determined for stars than for BDs.
+None the less, BDs that are wide companions to well-characterized stars provide ideal benchmarks [e.g., Gliese 229 B, (Nakajima et al., 1995) and Indi B, (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010)].
+Binary systems are expected to have coeval formation, sharing common age, metallicity, and distance – parameters more readily determined for stars than for BDs.
 This characteristic makes such systems particularly valuable for BD studies.
 The T dwarf spectral class was established to classify BDs cool enough to exhibit methane absorption in their near-infrared (NIR) spectra (Burgasser et al., 2002; Geballe et al., 2002).
-T dwarfs have effective temperatures () between 500–1300 K (Kirkpatrick et al., 2021).
+T dwarfs have effective temperatures () between 500 and 1300 K (Kirkpatrick et al., 2021).
 They constitute the majority of field BDs (e.g.
@@
 Best et al., 2021).
-Currently, there are about 30 T dwarfs known with stellar companions (summarised in Best et al., 2024), with significant contributions from the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010).
-To expand the parameter space of benchmark T dwarfs and improve their characterization, we searched for late-type T dwarfs with wide stellar companions using the CatWISE2020 catalog (Marocco et al., 2021) and the third data release (DR3) of Gaia (Gaia Collaboration et al., 2023).
-In this paper, we present the discovery of two new M + T dwarf wide binary systems.
+Currently, there are about 30 T dwarfs known with stellar companions (summarized in Best et al., 2024), with significant contributions from the Wide-field Infrared Survey Explorer (WISE; Wright et al., 2010).
+To expand the parameter space of benchmark T dwarfs and improve their characterization, we searched for late-type T dwarfs with wide stellar companions using the CatWISE2020 catalogue (Marocco et al., 2021) and the third data release (DR3) of Gaia (Gaia Collaboration et al., 2023).
+In this paper, we present the discovery of a wide binary of mildly metal-poor M2 + T5 dwarfs and a probable young hierarchical quadruple system of [M4 + M4] + [T7 + T8] dwarfs.
 Section 2 describes the selection of wide binary candidates.
@@
 2 Selection of wide binaries
-Wide binaries ( 100 au) exhibit negligible orbital motion over decadal timescales due to their extremely long orbital periods ( 1 kyr).
+Wide binaries ( 100 au) exhibit negligible orbital motion over decadal time-scales due to their extremely long orbital periods ( 1 kyr).
 In the solar neighbourhood ( 100 pc), the proper motions (PMs) of wide binary components significantly exceed their orbital motions, resulting in apparent common PM over decades.
@@
 T dwarfs emitting primarily in the infrared, only those in the solar neighbourhood are detectable by large-scale infrared surveys.
-The CatWISE database, providing mid-infrared photometry and PMs for nearby objects, serves as a powerful tool for identifying nearby T dwarf wide binaries.
-We focused on late-type T dwarfs, which exhibit redder colour (Eq.
-1) and suffer less contamination, e.g.
+The CatWISE catalogue, providing mid-infrared photometry and PMs for nearby objects, serves as a powerful tool for identifying nearby T dwarf wide binaries.
+We focused on late-type T dwarfs, which exhibit redder colour (equation 1) and suffer less contamination, e.g.
 from poorly measured background objects, than earlier types.
 Candidate selection from CatWISE employed the following criteria:
-In our selection, we required signal to noise ratio (SNR) of better than 5.0 for W1 and W2 detections (Eq.
-2).
+In our selection, we required signal-to-noise ratio (SNR) of better than 5.0 for and detections (equation 2).
 SNR of 5.0 corresponds to an error of 0.22 in W1 or W2 magnitudes.
-To mitigate contamination from background stars in the low-resolution (275 per pixel) unWISE images (Lang, 2014) used for CatWISE, we excluded the Galactic plane (Eq.
-5).
-We select objects without identified artifacts in both W1 and W2 band unWISE images (Eq.
-7) and with only one blend component used in the fit (Eq.
-6).
+To mitigate contamination from background stars in the low-resolution (2.75 arcsec pixel-1) unblurred coadds of the WISE images (unWISE; Lang, 2014) used for CatWISE, we excluded the Galactic plane (equation 5).
+We select objects without identified artefacts in both and band unWISE images (equation 7) and with only one blend component used in the fit (equation 6).
 Approximately 9,540 objects in CatWISE met these criteria (Eqs 1–7).
 We then searched for wide companions of these objects in the Gaia Catalogue of Nearby Stars (GCNS; Gaia Collaboration et al., 2021) using common PM. Our common PM pair criteria were:
-Given the relatively large PM errors in CatWISE, we required significant PM (Eq.
-3) but relaxed the PM constraint in the GCNS cross-match (Eq.
-8).
-The distance limit of 40 pc (Eq.
-9) corresponds to the -band detection limit of CatWISE for late T dwarfs (e.g.
+Given the relatively large PM errors in CatWISE, we required significant PM (equation 3) but relaxed the PM constraint in the GCNS cross-match (equation 8).
+The distance limit of 40 pc (equation 9) corresponds to the -band detection limit of CatWISE for late T dwarfs (e.g.
 table 6 of Zhang et al., 2019).
-We limited the binary separation to within 10′ (corresponding to 24,000 au at 40 pc), beyond which the number of random contaminations arises, partially due to the large PM errors in CatWISE.
+We limited the binary separation to within 10 arcmin (corresponding to 24,000 au at 40 pc), beyond which the number of random contaminations arises, partially due to the large PM errors in CatWISE.
 Our search yielded 15 candidate common PM pairs.
-After visual inspection using data from the Dark Energy Survey (DES; Abbott et al., 2021) and the DESI Legacy Imaging Surveys (Dey et al., 2019a), we excluded seven pairs of which the faint component has poor unWISE images, has Dark Energy Camera (DECam) deep -band non-detection (), or has no DECam observation and non-detection by other shallower surveys.
-The remaining eight T dwarf candidates, all with clear DECam -band detections, comprise two newly identified wide binaries and six known wide systems (Gliese 570 A & D, K4V + T7.5, Burgasser et al.
-2000; G 204-39 AB, M3+T6.5, Faherty et al.
-2010; SDSS J1416+13 AB, sdL7+sdT7.5, Scholz 2010; Burningham et al.
-2010; HIP 73786 AB, sdM4+sdT5.5, Murray et al.
-2011; Zhang et al.
-2019; LHS 6176 AB, K8V+T8p, Burningham et al.
-2013; L 34-26 + COCONUTS-2b, M3V+T9, Zhang et al.
-2021).
+After visual inspection using data from the Dark Energy Survey (Abbott et al., 2021) and the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Surveys (Dey et al., 2019), we excluded seven pairs of which the faint component has poor unWISE images, has Dark Energy Camera (DECam) deep -band non-detection (), or has no DECam observation and non-detection by other shallower surveys.
+The remaining eight T dwarf candidates, all with clear DECam -band detections, comprise two newly identified wide binaries and six known wide systems [Gliese 570 A & D, K4V + T7.5, (Burgasser et al., 2000); G 204-39 AB, M3+T6.5, (Faherty et al., 2010); SDSS J1416+13 AB, sdL7+sdT7.5, (Scholz, 2010; Burningham et al., 2010); HIP 73786 AB, sdM4+sdT5.5, (Murray et al., 2011; Zhang et al., 2019); LHS 6176 AB, K8V+T8p, (Burningham et al., 2013); L 34-26 + COCONUTS-2b, M3V+T9, (Zhang et al., 2021)].
 Fig.
@@
 Rothermich et al.
-(2024) classified UPM J1040–3551 A as an M3.8 dwarf based on photometric colours, and classified UPM J1040–3551 B as T7 dwarf based on a low resolution near infrared spectrum observed on 2023 June 30.
+(2024) classified UPM J1040–3551 A as an M3.8 dwarf based on photometric colours, and classified UPM J1040–3551 B as T7 dwarf based on a low-resolution NIR spectrum observed on 2023 June 30.
 Figure 2 shows the finding charts of these two new binaries.
@@
 L 122-88 AB and UPM J10403551 AB are located at distances of 33 and 25 pc respectively.
-They were missed by earlier searches with WISE and AllWISE (Wright et al., 2010; Mainzer et al., 2011), possibly because T dwarf candidates within 20 pc were prioritised in those spectroscopic follow up programs, to achieve a more complete volume-limited sample (e.g., Kirkpatrick et al., 2011, 2021; Mace et al., 2013).
-Another seven known T dwarf binaries were selected as T dwarf candidates by Eqs 1–7, but were not recovered as wide binaries mainly because of large uncertainties of their CatWISE PMs.
-We recovered most of them by double the limits in Eqs 3 and 4, but no new binary T dwarf candidate was found.
-Primary | L 122-88 A | UPM J10403551 A | Ref. |
+They were missed by earlier searches with WISE and AllWISE (Wright et al., 2010; Mainzer et al., 2011), possibly because T dwarf candidates within 20 pc were prioritized in those spectroscopic follow up programmes, to achieve a more complete volume-limited sample (e.g., Kirkpatrick et al., 2011, 2021; Mace et al., 2013).
+Another seven known T dwarf binaries were selected as T dwarf candidates by equations (1–7), but were not recovered as wide binaries mainly because of large uncertainties of their CatWISE PMs.
+We recovered most of them by double the limits in equations (3) and (4), but no new binary T dwarf candidate was found.
+Primary | L 122-88 A | UPM J10403551 Aa/Ab | Ref. |
 CWISE | J003124.30641354.9 | J104055.33355130.9 | |
@@
 (2016) | (1) | ||
-Spectral type | M2 | M4 | (2) |
+Spectral type | M2 | M4 + M4 | (2) |
 Age (Gyr) | 2.5-9.9 | 0.3-2.0 | (3,2) |
@@
 (2MASS) | 8.991 0.019 | 9.446 0.023 | (4) |
-W1 | 8.868 0.012 | 9.280 0.012 | (5) |
-W2 | 8.729 0.007 | 9.087 0.007 | (5) |
+8.868 0.012 | 9.280 0.012 | (5) | |
+8.729 0.007 | 9.087 0.007 | (5) | |
 (mas) | 30.206 0.013 | 39.552 0.020 | (1) |
@@
 (K) | 3500 50 | 3200 50 | (2) |
-Mass (M⊙) | 0.378 0.020 | 0.235 0.020 | (6) |
+Mass (M⊙) | 0.378 0.020 | 0.174 0.022/0.165 0.020 | (6,2) |
 Companion B | L 122-88 B | UPM J10403551 Ba/Bb | |
@@
 (WISE) | (3) | ||
-Mean mjd (day) | 57265.87 | 57124.73 | (3) |
+Mean MJD (d) | 57265.87 | 57124.73 | (3) |
 Spectral type | d/sdT5 | T7 + T8 | (2) |
@@
 (VHS) | 18.417 0.239 | 17.344 0.128 | (8) |
-W1 | 17.380 0.055 | 16.773 0.041 | (5) |
-W2 | 15.351 0.029 | 14.940 0.025 | (5) |
+17.380 0.055 | 16.773 0.041 | (5) | |
+15.351 0.029 | 14.940 0.025 | (5) | |
 Distance (pc) | 45 | a20, b29 | (2) |
@@
 Mass (M⊙) | 0.05–0.067 | 0.011–0.033/0.009–0.027 | (2) |
-Separation (″) | 215.6 | 65.48 | (2) |
+Separation (arcsec) | 215.6 | 65.48 | (2) |
 Proj. sep. (au) | 7138.4 0.5 | 1655.6 0.3 | (2) |
-cProj. sep. () | 0.034 | 0.009 | (2) |
+cProj. sep. () | 0.034 | 0.008 | (2) |
 (J) | (2) |
@@
 bAverage spectroscopic distance as a T7+T8 binary. The extracted apparent magnitude of the T7 companion is used in the calculation.
-cProjected separation in Jacobi radius (, see Section 4.3.)
-Name | SpT | UT date | Telescope | Instrument | Slit | Seeing | Airm | Wavelength | Resolution | |
-(arcsec) | (arcsec) | (nm) | () | (sec) | ||||||
+cProjected separation in Jacobi radius (, see Section 4.3).
+Name | SpT | UT date | Telescope | Instrument | Slit | Seeing | Airmass | Wavelength | Resolution | |
+(arcsec) | (arcsec) | (nm) | () | (s) | ||||||
 UPM J1040–3551 A | M5 | 2023-12-29 | SOAR | Goodman/400 | 1.0 | 1.45 | 1.04 | 492-897 | 2000 | |
@@
 As spectra for L 122-88 B and UPM J10403551 A were not available in the literature, we continued our spectroscopic campaign to complete the analysis of these new binary systems.
-Table 2 summarises the details of our observations.
+Table 2 summarizes the details of our observations.
 3.1 SOAR/Goodman
-L 122-88 A and UPM J1040–3551 A were observed using the Goodman High Throughput Spectrograph (Goodman; Clemens et al., 2004) at the Southern Astrophysical Research Telescope (SOAR) in long-slit mode on 2023 December 29 (see Fig.
+L 122-88 A and UPM J1040–3551 A were observed using the Goodman High Throughput Spectrograph (Goodman; Clemens et al., 2004) at the Southern Astrophysical Research (SOAR) telescope in long-slit mode on 2023 December 29 (see Fig.
 3).
-We used a slit and the 400 lines mm-1 grism in M2 mode, providing wavelength coverage from 492 to 897 nm with a resolving power of 2,000.
-Observations were conducted at airmasses of 1.89 and seeing of 1.55 for L 122-88 A, and airmasses of 1.04 and seeing of 1.45 for UPM J1040–3551 A.
-Data reduction was performed using the Goodman HST Pipeline111https://soardocs.readthedocs.io/projects/goodman-pipeline/ for basic CCD reduction, including bias subtraction, flat-field correction, and cosmic ray removal.
+We used a slit and the 400 lines mm-1 grism in M2 mode, providing wavelength coverage from 492 to 897 nm with a resolving power of 2000.
+Observations were conducted at airmass of 1.89 and seeing of 1.55 for L 122-88 A, and airmass of 1.04 and seeing of 1.45 for UPM J1040–3551 A.
+Data reduction was performed using the Goodman High Throughput Spectrograph Pipeline111https://soardocs.readthedocs.io/projects/goodman-pipeline/ for basic CCD reduction, including bias subtraction, flat-field correction, and cosmic ray removal.
 Further spectroscopic processing utilized tasks from the NOAO package in IRAF, including wavelength calibration, spectral extraction, telluric correction, and flux calibration.
@@
 UPM J10403551 B and L 122-88 B were observed with the TripleSpec4.1 NIR Imaging Spectrograph (TripleSpec; Schlawin et al., 2014) at SOAR.
-TripleSpec provides simultaneous spectral coverage from 940 to 2470 nm (R3,500), spanning the entire photometric range (see Fig.
+TripleSpec provides simultaneous spectral coverage from 940 to 2470 nm (3500), spanning the entire photometric range (see Fig.
 4).
-UPM J10403551 B was observed on 2024 March 26 (seeing: 091, airmass: 1.07), while L 122-88 B was observed on 2024 July 11 (seeing: 124, airmass: 1.35).
+UPM J10403551 B was observed on 2024 March 26 (seeing: 0.91 arcsec, airmass: 1.07), while L 122-88 B was observed on 2024 July 11 (seeing: 1.24 arcsec, airmass: 1.35).
 We adopted the default ABBA dither pattern with eight 225-s exposures for UPM J1040–3551 B and twenty 200-s exposures for L 122-88 B.
@@
 Telluric correction and flux calibration were performed using observations of the standard star HIP 54311.
-The resulting NIR spectra exhibit peak signal-to-noise ratios (SNR) of 10-15 at the centre of the -bands.
+The resulting NIR spectra exhibit peak SNRs of 10-15 at the centre of the bands.
 3.3 VLT/X-shooter
-The X-shooter spectrograph (Vernet et al., 2011) at the Very Large Telescope comprises three spectroscopic arms (UVB, VIS, and NIR), each optimized for its spectral range (300-2500 nm).
-The optical to NIR spectrum of L 122-88 A was obtained using the X-shooter on 2024 June 16 (seeing: 15, airmass: 1.8).
+The X-shooter spectrograph (Vernet et al., 2011) at the Very Large Telescope (VLT) comprises three spectroscopic arms [UV-Blue (UVB), Visible (VIS), and NIR], each optimized for its spectral range (300-2500 nm).
+The optical to NIR spectrum of L 122-88 A was obtained using the X-shooter on 2024 June 16 (seeing: 1.5 arcsec, airmass: 1.8).
 Figure 5 presents the combined, smoothed X-shooter spectrum of L 122-88 A.
-L 122-88 B was observed with X-shooter on 2024 July 7 (seeing: 358) and 2024 July 13 (seeing: 129).
-Figure 6 shows the X-shooter spectrum of L 122-88 B, noting that the optical spectrum is contaminated by a nearby galaxy, in fact, we note the presence of a non-zero continuum in the 400-800 nm spectral region in which T dwarfs should barely have emit any flux (e.g., Burgasser et al., 2003b).
-The nearby galaxy is faint in the -band (see Fig.
+L 122-88 B was observed with X-shooter on 2024 July 7 (seeing: 3.58 arcsec) and 2024 July 13 (seeing: 1.29 arcsec).
+Fig.
+6 shows the X-shooter spectrum of L 122-88 B; noting that the optical spectrum is contaminated by a nearby galaxy, in fact, we note the presence of a non-zero continuum in the 400-800 nm spectral region in which T dwarfs should barely have emitted any flux (e.g., Burgasser et al., 2003b).
+The nearby galaxy is faint in the band (see Fig.
 2) but brighter than L 122-88 B in the optical.
-With a PM of 0368 yr-1 towards the north-east, L 122-88 B has moved to the top of the galaxy in 2024 when the X-shooter spectrum was taken.
+With a PM of 0.368 arcsec yr-1 towards the north-east, L 122-88 B has moved to the top of the galaxy in 2024 when the X-shooter spectrum was taken.
 Meanwhile, the flux of the X-shooter spectrum beyond 2250 nm is poorly extracted/calibrated due to low SNR.
-These spectra were reduced using the ESO Reflex pipeline (Freudling et al., 2013).
+These spectra were reduced using the European Southern Observatory Reflex pipeline (Freudling et al., 2013).
 4 Characterization
 4.1 M dwarf companions
-The Gaia Renormalized Unit Weight Error (RUWE) values for L 122-88 A and UPM J10403551 A (1.32 and 1.47, respectively) suggest these objects may potentially be unresolved binaries.
+The Gaia Renormalized Unit Weight Error (RUWE) values for L 122-88 A and UPM J10403551 A (1.32 and 1.47, respectively) suggest that these objects may potentially be unresolved binaries.
 However, their optical spectra appear consistent with normal M dwarfs (Fig.
@@
 Under such circumstances, given their identical distances, metallicity constraints derived from model fitting, whether applied to their composite spectra or to individual components, would yield consistent results.
-In this section, we examine their spectra as single objects.
+In this section, we first examine their spectra as single objects, then discuss their binarity in Section 4.1.4.
 4.1.1 L 122-88 A
@@
 3) of L 122-88 A to those M0-M9 dwarf templates from Bochanski et al.
-(2007), and found it closely matches an M2 dwarf template.
-However, it exhibits slightly weaker TiO absorption-bands around 710 nm, which indicates a mildly metal-poor composition (e.g., Lépine et al., 2007; Jao et al., 2008; Zhang, 2019).
-Analysis of Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) data using the TESSILATOR software (Binks & Günther, 2024) reveals a rotation signal of 2.4750.013 days in sector 2 data.
+(2007), and found that it closely matches an M2 dwarf template.
+However, it exhibits slightly weaker TiO absorption bands around 710 nm, which indicates a mildly metal-poor composition (e.g., Lépine et al., 2007; Jao et al., 2008; Zhang, 2019).
+Analysis of data from the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2015) using the TESSILATOR software (Binks & Günther, 2024) reveals a rotation signal of 2.475 0.013 d in sector 2 data.
 Applying gyrochronology to this rotation period using GPgyro (Lu et al., 2024) yields an age estimate of 0.0290.010 Gyr.
-However, there is no evidence for this signal in other TESS sectors 28, 29, 68, and 69.
-Examination of the TESS data sector 2 release notes222https://ntrs.nasa.gov/citations/20190001284 for TIC 271503151 (L 122-88 A) does not suggest any particular issue with the sector 2 observation but for a V=13 object we would expect the signal to be visible in other sectors for a reliable detection.
+However, there is no evidence for this signal in other TESS sectors (28, 29, 68, and 69).
+Examination of the TESS data sector 2 release notes222https://ntrs.nasa.gov/citations/20190001284 for TIC 271503151 (L 122-88 A) does not suggest any particular issue with the sector 2 observation but for a = 13 object we would expect the signal to be visible in other sectors for a reliable detection.
 Marocco et al.
@@
 We compared the optical spectrum (Fig.
-3) of UPM J10403551 A to those M0-M9 dwarf templates from Bochanski et al.
+3) of UPM J10403551 A to those M0–M9 dwarf templates from Bochanski et al.
 (2007), and found it closely matches that of an M4 dwarf template.
@@
 We measured the H equivalent width (H EW) of UPM J10403551 A to be between 2.33 and 3.78 Å.
-Applying the H activity-age correlation for M3-6 dwarfs from Kiman et al.
+Applying the H activity–age correlation for M3–M6 dwarfs from Kiman et al.
 (2021) (their fig.
 6), this H EW range corresponds to an estimated age of 0.3-2.0 Gyr.
-The TESS data for UPM J10403551 A shows no indication of activity or any periodic signal.
+The TESS data for UPM J10403551 A show no indication of activity or any periodic signal.
 4.1.3 Atmospheric properties
 The optical spectrum is more sensitive to of M (sub)dwarfs, while the NIR spectrum is more sensitive to [Fe/H].
-Therefore, the full optical-NIR spectral profile of M subdwarfs can have better constraints of their and [Fe/H].
-To derive the atmospheric parameters of L 122-88 A, we fitted its X-shooter spectrum (300-2500 nm) to BT-Settl models with 2600 K K, [Fe/H] 0.5, and log = 5.0 (Allard, 2014).
-We used a typical log = 5.0 for M2-M4 dwarfs as log has relatively small affection on their spectra compared to and [Fe/H].
+Therefore, the full optical–NIR spectral profile of M subdwarfs can have better constraints of their and [Fe/H].
+To derive the atmospheric parameters of L 122-88 A, we fitted its X-shooter spectrum (300–2500 nm) to BT-Settl models with 2600 K K, [Fe/H] 0.5, and log = 5.0 (Allard, 2014).
+We used a typical log = 5.0 for M2–M4 dwarfs as log has relatively small effect on their spectra compared to and [Fe/H].
 We applied smoothing to the X-shooter spectrum to enhance the SNR, though this degraded the resolving power to 1000.
@@
 3), indicating a solar metallicity for this object.
+4.1.4 UPM J10403551 A: a likely near equal-mass binary
+Single stars are expected to have RUWE 1.0.
+Therefore, RUWE 1.4 is usually used as an indicator of non-single object.
+Nearby M dwarfs and earlier type stars predominantly exhibit RUWE values of 1.2 0.3 and 1.0 0.2, respectively (Sozzetti, 2023).
+The RUWE = 1.47 of UPM J10403551 A suggests it is likely an unresolved binary.
+For unresolved equal-mass binaries, the combined magnitude appears 0.75 mag brighter than each individual component, providing a method to test the binarity hypothesis.
+UPM J10403551 A exhibits a typical (3200 50 K) for M4 dwarfs (see fig.
+4 of Zhang et al., 2018a).
+Notably, model spectral fitting would yield identical temperature estimates regardless of whether the object is a single star or an equal-mass binary.
+According to evolutionary models (Baraffe et al., 2015), a solar-metallicity star with K and age of 0.5-5 Gyr should have a mass of 0.174 M⊙, radius of 0.195 R⊙, and absolute magnitudes of , , and .
+The observed absolute magnitudes of UPM J10403551 A (, , and ) are systematically brighter by 0.69, 0.73, and 0.70 mag, respectively, than predicted by models for its temperature (see Fig.
+7).
+This consistent 0.7 mag excess strongly supports the classification of UPM J10403551 A as a near equal-mass binary system (UPM J10403551 Aab), with components Aa and Ab differing by only 0.1 mag ( 25 K) and both can be classified as M4 dwarfs.
+As a single star, UPM J10403551 A would have a mass of M⊙ according to the mass- correlation (Mann et al., 2019).
+A solar-metallicity star with this mass and age of 0.5-5 Gyr should have a K, significantly higher than the model spectral fitting result.
+With of 3200 K and 3175 K, UPM J10403551 Aa & Ab should have a mass around 0.174 0.022 M⊙ and 0.165 0.020 M⊙, respectively, according to evolutionary models (Baraffe et al., 2015).
+For comparison, we also evaluated L 122-88 A by comparing its observed , , and values to evolutionary model predictions for a solar-metallicity star with K and age of 8 Gyr (see Fig.
+7).
+The resulting magnitude differences of approximately 0.11 mag are insufficient to suggest binarity in this case.
 4.2 T dwarf companions
@@
 The absence of the -band flux excess feature in L 122-88 B indicates a metallicity higher than that of T subdwarfs, which usually have [Fe/H] (e.g., fig.
-9, Zhang et al., 2017b).
+9 of Zhang et al., 2017b).
 This is still consistent with the metallicity derived from L 122-88 A.
-To further investigate the mildly metal-poor nature of L 122-88 B, we compared its absolute magnitudes in the and -bands, computed using the Gaia parallax of the primary, to those of known T dwarfs and subdwarfs (Fig.
-7).
+To further investigate the mildly metal-poor nature of L 122-88 B, we compared its absolute magnitudes in the and bands, computed using the Gaia parallax of the primary, to those of known T dwarfs and subdwarfs (Fig.
+8).
 L 122-88 B exhibits slightly fainter and compared to typical T5 dwarfs, while its remains similar.
@@
 Without considering its mildly metal-poor nature, L 122-88 B might be misclassified as a T6.5 1 dwarf based on spectral type versus absolute magnitude correlations (Fig.
-7).
+8).
 Indeed, Marocco et al.
@@
 We also estimated the distance of L 122-88 B by the spectral type – absolute magnitude correlations (fig.
-10, Zhang et al., 2019).
-The average distance estimated from its magnitudes is 45 pc, which is consistent to the Gaia distance of L 122-88 A (33.1060.014) within the uncertainty.
-Of course, L 122-88 B exhibits a over estimated spectral distance due to its mildly metal-poor composition.
+10 of Zhang et al., 2019).
+The average distance estimated from its magnitudes is 45 pc, which is consistent to the Gaia distance of L 122-88 A (33.106 0.014) within the uncertainty.
+Of course, L 122-88 B exhibits an over estimated spectral distance due to its mildly metal-poor composition.
 This metallicity deficiency causes its actual absolute magnitudes to be fainter than the average values for T5 dwarfs used in the calculation.
@@
 4 reveals flux excesses in the and bands relative to the T7 standard when normalized at the -band peak.
-When rescaling the spectrum of UPM J1040-3551 B to better align with the T7 standard across , , and bands, we observe a 13% flux suppression in the band, a phenomenon neither previously observed nor explained in any single BD.
-Several features of evidence suggest that UPM J10403551 B is not a T subdwarf: (1) it exhibits -band flux excess rather than the suppression typical of late-type T subdwarfs, (2) it has a relatively young age of 0.3-2.0 Gyr inferred from its primary (Section 4.1.2), and (3) it has solar metallicity as indicated by its primary’s spectrum (Fig.
+When rescaling the spectrum of UPM J10403551 B to better align with the T7 standard across , , and bands, we observe a 13% flux suppression in the band, a phenomenon neither previously observed nor explained in any single BD.
+Several features of evidence suggest that UPM J10403551 B is not a T subdwarf: (1) it exhibits -band flux excess rather than the suppression typical of late-type T subdwarfs; (2) it has a relatively young age of 0.3-2.0 Gyr inferred from its primary (Section 4.1.2); and (3) it has solar metallicity as indicated by its primary’s spectrum (Fig.
 3).
-While the -band flux excess could be attributed to its young age (e.g., Burningham et al., 2011), this explanation fails to account for the observed and band flux excesses.
-Unresolved binary BDs may show unusual spectral features and colours, and flux excesses if have been treated as single objects (Burgasser et al., 2010; Zhang et al., 2010; Marocco et al., 2015; Zhang et al., 2018b; Kirkpatrick et al., 2021).
-Figure 7 demonstrates that UPM J10403551 B has significantly brighter and band absolute magnitudes than typical T7 dwarfs.
-Such over-luminosity in field ultracool dwarfs often indicates the presence of an unresolved companion.
-To test this hypothesis, we conducted a binary spectral fitting analysis using single spectral templates in the Spex Prism Library Analysis Toolkit (Burgasser & Splat Development Team, 2017, SPLAT;).
-We selected 131 T4–T9 dwarfs with low-resolution ( 150) NIR spectra previously obtained with the SpeX spectrograph (Rayner et al., 2003) on the 3m NASA Infrared Telescope Facility.
+While the -band flux excess could be attributed to its young age (e.g., Burningham et al., 2011), this explanation fails to account for the observed - and -band flux excesses.
+Unresolved binary BDs may show unusual spectral features and colours, and flux excesses if they have been treated as single objects (Burgasser et al., 2010; Zhang et al., 2010; Marocco et al., 2015; Zhang et al., 2018b; Kirkpatrick et al., 2021).
+Figure 8 demonstrates that UPM J10403551 B has significantly brighter -, -, and -band absolute magnitudes than typical T7 dwarfs.
+Such overluminosity in field ultracool dwarfs often indicates the presence of an unresolved companion.
+To test this hypothesis, we conducted a binary spectral fitting analysis using single spectral templates in the Spex Prism Library Analysis Toolkit (SPLAT; Burgasser & Splat Development Team, 2017).
+We selected 131 T4–T9 dwarfs with low-resolution ( 150) NIR spectra previously obtained with the Spectral cross-disperser (SpeX) spectrograph (Rayner et al., 2003) on the 3m NASA Infrared Telescope Facility.
 These templates exclude known or suspected binaries and spectra with very low S/N 15.
@@
 We then co-added pairs of spectra to create 6,715 unique binary templates.
-We compared both single and binary templates to the equivalently interpolated SOAR spectrum of J10403551 B following procedures described in (Burgasser et al., 2010), masking out regions of strong telluric absorption between 1.35–1.42 m and 1.80–1.95 m.
-Figure 8 compares the best fit single template from this sample, the T7 dwarf WISE J004024.88+090054.8 (Mace et al., 2013, J0040+0900;) to the best-fit binary template composed of J0040+0900 and the T8 dwarf ULAS J102940.52+093514.6 (Burningham et al., 2013; Thompson et al., 2013, J1029+0935;).
+We compared both single and binary templates to the equivalently interpolated SOAR spectrum of J10403551 B following procedures described in Burgasser et al.
+(2010), masking out regions of strong telluric absorption between 1.35–1.42 m and 1.80–1.95 m.
+Figure 9 compares the best fitting single template from this sample, the T7 dwarf WISE J004024.88+090054.8 (Mace et al., 2013, J0040+0900;) to the best-fitting binary template composed of J0040+0900 and the T8 dwarf ULAS J102940.52+093514.6 (J1029+0935; Burningham et al., 2013; Thompson et al., 2013).
 The latter is a superior fit, notably providing a better reproduction of the , , and -band peaks.
@@
 Using the spectral type versus absolute magnitude correlations for L and T dwarfs (Fig.
-7), we derived , , and differences between T7 and T8 dwarfs of 1.258, 1.248, 0.907, and 0.462 mag, respectively.
-The decomposed , , and values demonstrate improved alignment with these correlations, though they remain slightly brighter than average—contrary to the dimmer characteristics of late T subdwarfs.
-UPM J10403551 B appears underluminous as an unresolved binary only in the -band, which may be attributed to its young age.
-For instance, the young T5.5 dwarf SDSSp J111010.01+011613.1 (Geballe et al., 2002) is relatively faint in the -band, with a colour 0.34 mag bluer than the average for field T5.5 dwarfs (fig.
-8i of Primeval VI, ).
-The independent spectroscopic distance of UPM J10403551 B as a single T7 dwarf is 20 pc, which is 21% shorter than the Gaia distance of the M4 primary (25.2830.013 pc).
-We also decomposed the band apparent magnitudes of UPM J10403551 B as T7 + T8 close binary.
-The average spectroscopic distance of UPM J10403551 Ba as a T7 dwarf is 21 pc, which is 17% smaller than its UPM J10403551 A but consistent within the uncertainty.
+8), we derived , , and differences between T7 and T8 dwarfs of 1.258, 1.248, 0.907, and 0.462 mag, respectively.
+The decomposed , , and values demonstrate improved alignment with these correlations, though they remain slightly brighter than average – contrary to the dimmer characteristics of late T subdwarfs.
+UPM J10403551 B appears underluminous as an unresolved binary only in the band, which may be attributed to its young age.
+For instance, the young T5.5 dwarf SDSSp J111010.01+011613.1 (Geballe et al., 2002) is relatively faint in the band, with a colour 0.34 mag bluer than the average for field T5.5 dwarfs (fig.
+8i of Zhang et al., 2019).
+The independent spectroscopic distance of UPM J10403551 B as a single T7 dwarf is 20 pc, which is 21 percent shorter than the Gaia distance of the M4 primary (25.2830.013 pc).
+We also decomposed the -, -, -, -band apparent magnitudes of UPM J10403551 B as T7 + T8 close binary.
+The average spectroscopic distance of UPM J10403551 Ba as a T7 dwarf is 21 pc, which is 17 percent smaller than its UPM J10403551 A but consistent within the uncertainty.
 The spectral distance of UPM J10403551 B is under estimated due to its young age, which allows it to have brighter absolute magnitudes than the average values for T7 dwarfs used in the calculation.
 Both the spectral fitting and the decomposition of absolute magnitudes strongly suggest that UPM J10403551 B is likely an unresolved T7 + T8 binary (UPM J10403551 Ba & Bb).
-UPM J10403551 ABab might be the third known triple system comprising a close binary of T dwarfs with a wide stellar companion, similar to Indi ABab (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010) or Gliese 229 ABab (Xuan et al., 2024).
-UPM J10403551 B is detected but not resolved in the Visible and Infrared Survey Telescope for Astronomy (VISTA) Hemisphere Survey (VHS; McMahon et al., 2013) and the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019b).
-The VHS and DECaLS images have pixel sizes of 034 and 026, respectively.
-This allows us to place upper limits on the orbital semi-major axis ( au) and period ( yr) of the system.
-At a distance of 25 pc, the components can potentially can be resolved at 1.15 m by the Near-Infrared Camera (NIRCam) of the James Webb Space Telescope (JWST), provided the orbital period exceeds six years.
+UPM J10403551 AabBab is the first known quadruple system comprising a close binary of T dwarfs in a wide orbit around a close binary of stars.
+It is the third identified system featuring a close binary of T dwarfs with wide stellar companion, following Indi ABab (Scholz et al., 2003; McCaughrean et al., 2004; King et al., 2010) and Gliese 229 ABab (Xuan et al., 2024).
+UPM J10403551 B is detected but not resolved in the Visible and Infrared Survey Telescope for Astronomy (VISTA) Hemisphere Survey (VHS; McMahon et al., 2013) and the Dark Energy Camera Legacy Survey (DECaLS; Dey et al., 2019).
+The VHS and DECaLS images have pixel sizes of 0.34 and 0.26 arcsec, respectively.
+This allows us to place upper limits on the orbital semimajor axis ( au) and period ( yr) of the system.
+At a distance of 25 pc, the components can potentially be resolved at 1.15 m by the Near-Infrared Camera of the JWST, provided the orbital period exceeds six years.
 4.3 Binary properties
@@
 As illustrated in Fig.
-9, L 122-88 B has an estimated mass of 0.05–0.067 M for an age of 2.5-9.9 Gyr, while UPM J10403551 Ba and Bb have estimated masses of 0.012–0.04 M and 0.007–0.02 M⊙, respectively, for an age of 0.3-2.0 Gyr.
-Using the derived distances, projected separations, and masses of all components, we calculated the binding energies of L 122-88 AB and UPM J10403551 ABab to be J and J, respectively.
+10, L 122-88 B has an estimated mass of 0.05–0.067 M for an age of 2.5-9.9 Gyr, while UPM J10403551 Ba and Bb have estimated masses of 0.012–0.04 M and 0.007–0.02 M⊙, respectively, for an age of 0.3-2.0 Gyr.
+Using the derived distances, projected separations, and masses of all components, we calculated the binding energies of L 122-88 AB and UPM J10403551 AabBab to be and J, respectively.
 To assess the stability of these wide binary systems, we considered the Jacobi (or tidal) radius (), which represents the maximum separation for stable wide binary systems.
@@
 For wide binaries in the solar neighbourhood, can be calculated using equation (43) from Jiang & Tremaine (2010).
-We found that L 122-88 AB and UPM J10403551 ABab have values of 1.02 pc and 0.87 pc, respectively.
-Their projected separations correspond to 0.034 and 0.009 , respectively, which are far smaller than their respective values.
+We found that L 122-88 AB and UPM J10403551 AabBab have values of 1.02 and 0.97 pc, respectively.
+Their projected separations correspond to 0.034 and 0.008 , respectively, which are far smaller than their respective values.
 This analysis strongly suggests that both systems are gravitationally stable.
@@
 5 Summary and conclusions
-We have conducted a search for late T dwarf wide binaries using CatWISE and Gaia DR3, resulting in the discovery of two new M + T dwarf wide binary systems.
+We have conducted a search for late T dwarf wide binaries using CatWISE and Gaia DR3, resulting in the discovery of a wide binary and a probable hierarchical quadruple system.
 Both systems are confirmed to be gravitationally bound and stable based on their tidal radii.
 Optical and NIR spectroscopic follow-up observations were performed for all companions, enabling spectral classification and characterization.
-The first system, L 122-88 AB, exhibits spectral features indicative of mild metal-poor composition, including a weaker 710 nm TiO absorption band in L 122-88 A and suppressed K-band flux in L 122-88 B.
+The first system, L 122-88 AB, exhibits spectral features indicative of mildly metal-poor composition, including a weaker 710 nm TiO absorption band in L 122-88 A and suppressed -band flux in L 122-88 B.
 The best-fitting BT-Settl model to the 0.3-2.5 m spectrum of L 122-88 A yields = 3500 K, [Fe/H] = , and log = 5.0.
 We classify L 122-88 AB as a mildly metal-poor M2 + T5 wide binary.
-The second system, UPM J10403551 AB, is likely a hierarchical triple system consisting of an unresolved T7 + T8 binary and an M4 dwarf wide companion at a distance of 25 pc from the Sun.
+The second system, UPM J10403551 AB, is likely a hierarchical quadruple system consisting of a T7 + T8 spectral binary and an M4 + M4 astrometric binary at a distance of 25 pc from the Sun.
+Both the RUWE (1.47) of UPM J10403551 A and its excess luminosity (0.7 mag brighter than model predications of its ) suggest it is an unresolved M4 + M4 binary.
 While UPM J10403551 Ba/Bb are unresolved in survey images, their NIR spectrum is best fitted by a combined T7 + T8 template, explaining the observed excess luminosity.
-This system might be the third known triple system similar to Indi ABab or Gliese 229 ABab.
+UPM J10403551 AabBab is the first known quadruple system with a close binary of T dwarfs.
 Based on the H emission of the M4 primary, we estimate the system age to be 0.3-2 Gyr.
@@
 Acknowledgements
-Based on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU).
-Based on observations collected at the European Southern Observatory under ESO programme 113.26V7.001.
+This study is based on observations obtained at the SOAR telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU).
+This study is based on observations collected at the European Southern Observatory under ESO programme 113.26V7.001.
 This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium).
 Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
-This publication makes use of data products from the Wide field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Lab- oratory/California Institute of Technology, funded by the National Aeronautics and Space Administration This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration.
-SPLAT is an collaborative project of research students in the UCSD Cool Star Lab, aimed at developing research through the building of spectral analysis tools.
+This publication makes use of data products from the WISE, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration.
+SPLAT is a collaborative project of research students in the UCSD Cool Star Lab, aimed at developing research through the building of spectral analysis tools.
 Contributors to SPLAT have included Christian Aganze, Jessica Birky, Daniella Bardalez Gagliuffi, Adam Burgasser (PI), Caleb Choban, Andrew Davis, Ivanna Escala, Joshua Hazlett, Carolina Herrara Hernandez, Elizabeth Moreno Hilario, Aishwarya Iyer, Yuhui Jin, Mike Lopez, Dorsa Majidi, Diego Octavio Talavera Maya, Alex Mendez, Gretel Mercado, Niana Mohammed, Johnny Parra, Maitrayee Sahi, Adrian Suarez, Melisa Tallis, Tomoki Tamiya, Chris Theissen, and Russell van Linge.
-Z.H.Z acknowledges the supports from the Jiangsu Province fundamental research programme (BK20211143), the Program for Innovative Talents and Entrepreneur in Jiangsu (JSSCTD202139), and the science research grants from the China Manned Space Project with NO.
+ZHZ acknowledges the supports from the Jiangsu Province fundamental research programme (BK20211143), the Program for Innovative Talents and Entrepreneur in Jiangsu (JSSCTD202139), and the science research grants from the China Manned Space Project with no.
 CMS-CSST-2021-A08.
-The work of F.N.
-is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
-M.C.G.O.
-acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF “A way of making Europe” through project PID2022-137241NB-C42.
-N.L.
-acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (AEI-MCINN) under grant PID2022-137241NB-C41.
-B.G.
-acknowledges support from the Polish National Science Center (NCN) under SONATA grant No.
+The work of FN is supported by NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation.
+MCGO acknowledges financial support from the Agencia Estatal de Investigación (AEI/10.13039/501100011033) of the Ministerio de Ciencia e Innovación and the ERDF ’A way of making Europe’ through project PID2022-137241NB-C42.
+NL acknowledges support from the Agencia Estatal de Investigación del Ministerio de Ciencia e Innovación (AEI-MCINN) under grant PID2022-137241NB-C41.
+BG acknowledges support from the Polish National Science Center (NCN) under SONATA grant no.
 2021/43/D/ST9/0194.
-P.C.
-acknowledges financial support from the Spanish Virtual Observatory project funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 through grant PID2020-112949GB-I00.
+PC acknowledges financial support from the Spanish Virtual Observatory project funded by the Spanish Ministry of Science and Innovation/State Agency of Research MCIN/AEI/10.13039/501100011033 through grant PID2020-112949GB-I00.
 The authors would like to thank the anonymous reviewer for his/her valuable comments.